# Microwave-to-Optical Quantum Transduction with Antiferromagnets

### v1

@@ I Introduction
Antiferromagnets have attract attention in the field of spintronics [1]. Unlike ferromagnets, antiferromagnets do not generate unwanted stray fields because of the zero net magnetization, which can be a merit of utilizing antiferromagnets near the magnetic-field sensitive devices such as superconducting qubits. Moreover, since the dynamics of antiferromagnets such as the antiferromagnetic resonance is typically in the terahertz regime, fast manipulation and utilization of antiferromagnet-based devices are expected. For example, it has been suggested that antiferromagnets can be active elements of a memory [2] and logic device [3], complementing or replacing ferromagnets.
@@
This paper is organized as follows. In Sec. II we introduce a generic lattice model for antiferromagnets with two sublattices and express it in terms of the antiferromagnetic magnon operators. In Sec. III we derive the interaction between microwave cavity photons and antiferromagnetic magnons from the Zeeman interaction in antiferromagnets. In Sec. IV we derive the interaction between optical cavity photons and antiferromagnetic magnons in the case with an optical cavity, and the interaction between optical itinerant photons and antiferromagnetic magnons in the case without an optical cavity, both from the magneto-optic Faraday effect in antiferromagnets. In Sec. V, starting from a generic description for calculating the transduction efficiency with the input-output formalism, we derive analytical expressions for the transduction efficiency of the microwave-to-optical quantum transduction in the cases with and without an optical cavity. In Sec. VI we present the numerical results for the transduction efficiency. In Sec. VII we briefly discuss a possible experimental realization and the merits of our proposal utilizing antiferromagnets. We also discuss possible ways to improve the transduction efficiency. In Sec. VIII we summarize this study.
@@ II Theoretical Model
To quadratic order in the sublattice magnon operators $\hat{a}_{\bm{k}}$ and $\hat{b}_{\bm{k}}$ in the Holstein-Primakoff transformation [37] which satisfy the commutation relations $[\hat{a}_{\bm{k}},\hat{a}^{\dagger}_{\bm{k}^{\prime}}]=\delta_{\bm{k},\bm{k}^{\prime}}$, $[\hat{b}_{\bm{k}},\hat{b}^{\dagger}_{\bm{k}^{\prime}}]=\delta_{\bm{k},\bm{k}^{\prime}}$, $[\hat{a}_{\bm{k}},\hat{a}_{\bm{k}^{\prime}}]=0$, and $[\hat{b}_{\bm{k}},\hat{b}_{\bm{k}^{\prime}}]=0$, it follows that
@@
where the matrix elements in Eq. (3) are determined under the condition such that the operators $\hat{m}_{\alpha,\bm{k}}$ and $\hat{m}_{\beta,\bm{k}}$ must satisfy the relations $[\hat{m}_{\alpha,\bm{k}},H_{\mathrm{AFM}}]=\hbar\omega_{\alpha,\bm{k}}\hat{m}_{\alpha,\bm{k}}$ and $[\hat{m}_{\beta,\bm{k}},H_{\mathrm{AFM}}]=\hbar\omega_{\beta,\bm{k}}\hat{m}_{\beta,\bm{k}}$.
@@
where $\alpha$ and $\beta$ denote the two antiferromagnetic magnon modes and $\bm{k}$ is the wave vector of the magnons. Henceforth we focus on the antiferromagnetic resonance state where all the spins are precessing uniformly (i.e., with $\bm{k}=0$) and drop the wave-vector dependence in $\omega_{\mu,\bm{k}}$ and $\hat{m}_{\mu,\bm{k}}$. The antiferromagnetic resonance frequencies are given by $\omega_{\alpha,\beta}^{2}=\omega_{E}(2\omega_{\parallel}+\omega_{\perp})+\omega_{H}^{2}\pm\sqrt{\omega_{E}^{2}\omega_{\perp}^{2}+4\omega_{H}^{2}\omega_{E}(2\omega_{\parallel}+\omega_{\perp})}$ with $\omega_{\alpha}\geq\omega_{\beta}$ [38, 34]. Here, the frequencies are given by $\omega_{E}=JSZ/\hbar$, $\omega_{H}=|\gamma|B_{0}$, $\omega_{\parallel}=SK_{\parallel}/\hbar$, and $\omega_{\perp}=SK_{\perp}/\hbar$, with $Z$ being the number of nearest neighbors in the antiferromagnet. A schematic illustration of the magnetic-field dependence of the resonance frequencies $\omega_{\alpha}$ and $\omega_{\beta}$ of easy-axis antiferromagnets is shown in Fig. 1(b).
@@ III Microwave-Magnon Interaction in Antiferromagnets
$\delta m_{x}$ $=\sqrt{S/2N}(U+V)\left(\alpha+\alpha^{\dagger}+\beta+\beta^{\dagger}\right),$
$\delta m_{y}$ $=(1/i)\sqrt{S/2N}(U+V)\left(\alpha-\alpha^{\dagger}-\beta+\beta^{\dagger}\right)$ (7)
@@
with $g_{0}=\eta|\gamma|\sqrt{\hbar\omega_{\mathrm{e}}\mu_{0}/(4V_{\mathrm{c}})}$. Here, $\eta\leq 1$ is the spatial overlap factor between the cavity mode and the antiferromagnet and $V_{\mathrm{c}}$ is the volume of the cavity. We note that only the expression for the coupling strength for the lower-energy mode $g_{\beta}$ in Eq. (10) was obtained in Ref. [34].
@@ IV Light-Magnon Interaction in Antiferromagnets
In this section, we derive the light-magnon interaction in antiferromagnets in the cases with and without an optical cavity, both from the magneto-optic Faraday effect in antiferromagnets. For concreteness, as in Sec. III, we focus on the easy-axis antiferromagnets with $K_{\perp}=0$.
@@ IV.1 The case with an optical cavity
where $\hat{b}$ is the annihilation operator of a photon with a (right or left) circular polarization. If we assume the equivalent Faraday rotation angles for the two sublattices $A$ and $B$, the coupling strength $G_{\mu}$ ($\mu=\alpha,\beta$) is given by [36]
$G_{\mu}=\frac{\kappa_{\mu}}{\sqrt{2SN}}\frac{c\theta_{\mathrm{F}}}{4\sqrt{\varepsilon}},$ (12)
where $\kappa_{\mu}$ is determined from the coefficients in the Bogoliubov transformation (3), $c$ is the speed of light, $\theta_{\mathrm{F}}$ is the Faraday rotation angle per unit length, and $\varepsilon$ is the relative permittivity of the antiferromagnet. In the case of easy-axis antiferromagnets with $\omega_{\parallel}/\omega_{E}\ll 1$, we have
@@
where $K$ denotes an intrinsic magneto-optic asymmetry between two sublattices [36]. Note that the expression for $\kappa_{\mu}$ in Eq. (13) is independent of the external magnetic field.
@@
$H_{\zeta}=-\hbar\left[\zeta_{\alpha}\left(\hat{m}_{\alpha}^{\dagger}+\hat{m}_{\alpha}\right)+\zeta_{\beta}\left(\hat{m}_{\beta}^{\dagger}+\hat{m}_{\beta}\right)\right]\left(\delta\hat{b}^{\dagger}+\delta\hat{b}\right),$ (14)
@@ IV.2 The case without an optical cavity
where $\hat{b}_{R,L}$ is the annihilation operator of a photon with a right-circular (left-circular) polarization. For concreteness, let us consider an incident light propagating in the $x$ direction (which is a direction perpendicular to the ground-state direction of the magnetic moments). For a strong $z$-polarized light we have $\hat{b}_{R,L}(t)=(\hat{b}_{y}\mp i\hat{b}_{z})/\sqrt{2}\simeq(\hat{b}_{y}\mp i\langle\hat{b}_{z}\rangle)/\sqrt{2}$, where $\langle\hat{b}_{z}\rangle=\sqrt{P_{0}/(\hbar\Omega_{0})}e^{-i\Omega_{0}t}$ with $P_{0}$ ($\Omega_{0}$) the power (angular frequency) of the incident light [41]. Because in the present case $\hat{b}$ and $\hat{b}^{\dagger}$ represent itinerant photon operators whose dimension is $[\mathrm{time}]^{-1/2}$, the interaction Hamiltonian (16) needs to be integrated over the interaction time [30].
[…]
@@
×$\times\left(\hat{b}_{\mathrm{in}}e^{i\Omega_{0}t}-\hat{b}^{\dagger}_{\mathrm{in}}e^{-i\Omega_{0}t}\right),$ (17)
@@
with $\phi_{\mathrm{F}}=\theta_{\mathrm{F}}d_{\mathrm{AFI}}$ being the total Faraday rotation angle of the antiferromagnet (which should not be confused with $\theta_{\mathrm{F}}$).
@@ V Microwave-to-Optical Quantum Transduction
where $H_{0}=\hbar\omega_{\mathrm{e}}\hat{a}^{\dagger}\hat{a}+\hbar\delta\omega_{\mathrm{o}}\hat{b}^{\dagger}\hat{b}+\hbar\omega_{\alpha}\hat{m}_{\alpha}^{\dagger}\hat{m}_{\alpha}+\hbar\omega_{\beta}\hat{m}_{\beta}^{\dagger}\hat{m}_{\beta}$ for systems with microwave and optical cavities, while $H_{0}=\hbar\omega_{\mathrm{e}}\hat{a}^{\dagger}\hat{a}+\hbar\omega_{\alpha}\hat{m}_{\alpha}^{\dagger}\hat{m}_{\alpha}+\hbar\omega_{\beta}\hat{m}_{\beta}^{\dagger}\hat{m}_{\beta}$ for systems with a microwave cavity but without an optical cavity. Here, $\delta\omega_{\mathrm{o}}\equiv\omega_{\mathrm{P}}-\omega_{\mathrm{o}}$ is the detuning of the optical cavity frequency from the pump frequency [29].
@@ V.1 General consideration
[…]
@@ V.2 The case with an optical cavity
$\dot{\hat{b}}=-i\delta\omega_{\mathrm{o}}\hat{b}-i\sum_{\mu}\zeta_{\mu}\hat{m}_{\mu}-\frac{\kappa_{\mathrm{o}}}{2}\hat{b}-\sqrt{\kappa_{\mathrm{o,e}}}\hat{b}_{\mathrm{in}},$ (26)
@@
$A$ $=\begin{bmatrix}i\omega_{\mathrm{e}}+\kappa_{\mathrm{e}}/2&ig_{\alpha}&ig_{\beta}&0\\ ig_{\alpha}&i\omega_{\alpha}+\gamma_{\alpha}/2&0&i\zeta_{\alpha}\\ ig_{\beta}&0&i\omega_{\beta}+\gamma_{\beta}/2&i\zeta_{\beta}\\ 0&i\zeta_{\alpha}&i\zeta_{\beta}&i\delta\omega_{\mathrm{o}}+\kappa_{\mathrm{o}}/2\end{bmatrix}$ (28)
@@
respectively. Finally, introducing the susceptibilities $\chi_{\mathrm{e}}=[-i(\omega-\omega_{\mathrm{e}})+\kappa_{\mathrm{e}}/2]^{-1}$, $\chi_{\mu}=[-i(\omega-\omega_{\mu})+\kappa_{\mu}/2]^{-1}$, and $\chi_{\mathrm{o}}=[-i(\omega-\delta\omega_{\mathrm{o}})+\kappa_{\mathrm{o}}/2]^{-1}$, the transduction efficiency $\eta$ of the microwave-to-optical quantum transduction is obtained from Eq. (23) as

### v2

@@ I Introduction
Antiferromagnets have attracted attention in the field of spintronics [1]. Unlike ferromagnets, antiferromagnets do not generate unwanted stray fields because of the zero net magnetization, which can be a merit of utilizing antiferromagnets near the magnetic-field sensitive devices such as superconducting qubits. Moreover, since the dynamics of antiferromagnets such as the antiferromagnetic resonance is typically in the terahertz regime, fast manipulation and utilization of antiferromagnet-based devices are expected. For example, it has been suggested that antiferromagnets can be active elements of a memory [2] and logic device [3], complementing or replacing ferromagnets.
@@
This paper is organized as follows. In Sec. II we introduce a generic lattice model for antiferromagnets with two sublattices and express it in terms of the antiferromagnetic magnon operators. In Sec. III we derive the interaction between microwave cavity photons and antiferromagnetic magnons from the Zeeman interaction in antiferromagnets. In Sec. IV we derive the interaction between optical cavity photons and antiferromagnetic magnons in the case with an optical cavity, and the interaction between optical itinerant photons and antiferromagnetic magnons in the case without an optical cavity, both from a linear magneto-optic (Faraday) effect in antiferromagnets. In Sec. V, starting from a generic description for calculating the transduction efficiency with the input-output formalism, we derive analytical expressions for the transduction efficiency of the microwave-to-optical quantum transduction in the cases with and without an optical cavity. In Sec. VI we present the numerical results for the transduction efficiency. In Sec. VII we briefly discuss a possible experimental realization and the merits of our proposal utilizing antiferromagnets. We also discuss possible ways to improve the transduction efficiency. In Sec. VIII we summarize this study.
@@ II Theoretical Model
To quadratic order in the sublattice magnon operators $\hat{a}_{\bm{k}}$ and $\hat{b}_{\bm{k}}$ in the Holstein-Primakoff transformation [37] which satisfy the commutation relations $\hat{a}_{\bm{k}},\hat{a}^{\dagger}_{\bm{k}^{\prime}}]=\delta_{\bm{k},\bm{k}^{\prime}}$, $\hat{b}_{\bm{k}},\hat{b}^{\dagger}_{\bm{k}^{\prime}}]=\delta_{\bm{k},\bm{k}^{\prime}}$, $\hat{a}_{\bm{k}},\hat{a}_{\bm{k}^{\prime}}]=0$, and $\hat{b}_{\bm{k}},\hat{b}_{\bm{k}^{\prime}}]=0$, it follows that
@@
where the matrix elements in Eqs. (3) and (4) are determined under the condition such that the operators $\hat{m}_{\alpha,\bm{k}}$ and $\hat{m}_{\beta,\bm{k}}$ must satisfy the relations $\hat{m}_{\alpha,\bm{k}},H_{\mathrm{AFM}}]=\hbar\omega_{\alpha,\bm{k}}\hat{m}_{\alpha,\bm{k}}$ and $\hat{m}_{\beta,\bm{k}},H_{\mathrm{AFM}}]=\hbar\omega_{\beta,\bm{k}}\hat{m}_{\beta,\bm{k}}$.
@@
where $\alpha$ and $\beta$ denote the two antiferromagnetic magnon modes and $\bm{k}$ is the wave vector of the magnons. Henceforth we focus on the antiferromagnetic resonance state where all the spins are precessing uniformly (i.e., with $\bm{k}=0$) and drop the wave-vector dependence in $\omega_{\mu,\bm{k}}$ and $\hat{m}_{\mu,\bm{k}}$. The antiferromagnetic resonance frequencies are given by $\omega_{\alpha,\beta}^{2}=\omega_{E}(2\omega_{\parallel}+\omega_{\perp})+\omega_{H}^{2}\pm\sqrt{\omega_{E}^{2}\omega_{\perp}^{2}+4\omega_{H}^{2}\omega_{E}(2\omega_{\parallel}+\omega_{\perp})}$ with $\omega_{\alpha}\geq\omega_{\beta}$ [38, 34]. Here, the frequencies are given by $\omega_{E}=SZJ/\hbar$, $\omega_{H}=|\gamma|B_{0}$, $\omega_{\parallel}=SK_{\parallel}/\hbar$, and $\omega_{\perp}=SK_{\perp}/\hbar$, with $Z$ being the number of nearest neighbors in the antiferromagnet. A schematic illustration of the magnetic-field dependence of the resonance frequencies $\omega_{\alpha}$ and $\omega_{\beta}$ of easy-axis antiferromagnets is shown in Fig. 1(b).
@@ III Microwave-Magnon Interaction in Antiferromagnets
$\delta m_{x}$ $=\sqrt{S/2N}(U+V)\left(\hat{m}_{\alpha}+\hat{m}_{\alpha}^{\dagger}+\hat{m}_{\beta}+\hat{m}_{\beta}^{\dagger}\right),$
$\delta m_{y}$ $=(1/i)\sqrt{S/2N}(U+V)\left(\hat{m}_{\alpha}-\hat{m}_{\alpha}^{\dagger}-\hat{m}_{\beta}+\hat{m}_{\beta}^{\dagger}\right)$ (7)
@@
with $g_{0}=\eta|\gamma|\sqrt{\hbar\omega_{\mathrm{e}}\mu_{0}/(4V_{\mathrm{c}})}$. Here, $\eta\leq 1$ is the spatial overlap factor between the cavity mode and the antiferromagnet and $V_{\mathrm{c}}$ is the volume of the cavity. We note that only the expression for the coupling strength for the lower-energy mode $g_{\beta}$ in Eq. (10) was obtained in Ref. [34]. In what follows, we drop the $\lambda$ dependence in Eq. (9) for simplicity.
@@ IV Light-Magnon Interaction in Antiferromagnets
In this section, we derive the light-magnon interaction in antiferromagnets in the cases with and without an optical cavity, both from a linear magneto-optic (Faraday) effect in antiferromagnets. See Appendix A for a brief derivation and for a comparison with the the light-magnon interaction in ferromagnets. For concreteness, as in Sec. III, we focus on the easy-axis antiferromagnets with $K_{\perp}=0$.
@@ IV.1 The case with an optical cavity
where $\hat{b}$ is the annihilation operator of a photon with a (right or left) circular polarization. If we assume an equal (absolute) value of the Faraday rotation angles for the two sublattices $A$ and $B$, the coupling strength $G_{\mu}$ ($\mu=\alpha,\beta$) is given by [36]
$G_{\mu}=\frac{c\theta_{\mathrm{F}}}{4\sqrt{\varepsilon_{\mathrm{r}}}}\frac{\kappa_{\mu}}{\sqrt{2SN}},$ (12)
where $\kappa_{\mu}$ is determined from the coefficients in the Bogoliubov transformation (3), $c$ is the speed of light, $\theta_{\mathrm{F}}$ is the Faraday rotation angle of each sublattice per unit length, and $\varepsilon_{\mathrm{r}}$ is the relative permittivity of the antiferromagnet. In the case of easy-axis antiferromagnets where the condition $\omega_{\parallel}/\omega_{E}\ll 1$ is usually satisfied, we have
@@
where $K=K_{-}/K_{+}$ denotes an intrinsic magneto-optic asymmetry between two sublattices (see Appendix A) [36]. Note that the expression for $\kappa_{\mu}$ in Eq. (13) is independent of the external magnetic field.
@@
[…]
@@
In the present context, the meaning of the Faraday effect in antiferromagnets may be unclear. Generically, the antiferromagnetic magnons can couple to photons through one-magnon and two-magnon Raman scattering processes. It is known that in antiferromagnets the magnitude of the one-magnon Raman scattering is rather small compared to that of the two-magnon Raman scattering [41, 42]. However, we here would like to emphasize that the light-magnon interaction in our study, which corresponds to a one-magnon scattering process, can be significantly enhanced by the presence of an optical cavity with the factor $\sqrt{\bar{n}_{\mathrm{cav}}}\sim 10^{3}$ [see Eq. (15)]. This one-magnon scattering, characterized by the factor $\theta_{\mathrm{F}}$, describes the Faraday rotation per sublattice [36] (see also Appendix A), which means that the total Faraday rotation angle can be zero at zero magnetic field in simple two-sublattice systems as is intuitively understood. On the one hand, it has also been suggested that in some antiferromagnets under magnetic fields the linear magneto-optic coupling (i.e., the Faraday rotation angle) can be as large as that of the ferromagnet YIG [43]. Due to the lack of the experimental data on the value of the Faraday rotation angle per sublattice in antiferromagnets, in Sec. VI we shall firstly compute the magnitude of the transduction efficiency as a function of $\theta_{\mathrm{F}}$. After that, for concreteness we shall adopt the value for YIG as an estimate of $\theta_{\mathrm{F}}$.
@@ IV.2 The case without an optical cavity
where $\hat{b}_{R,L}$ is the annihilation operator of a photon with a right-circular (left-circular) polarization. For concreteness, let us consider an incident light propagating in the $x$ direction (which is a direction perpendicular to the ground-state direction of the magnetic moments). For a strong $z$-polarized light we have $\hat{b}_{R,L}(t)=(\hat{b}_{y}\mp i\hat{b}_{z})/\sqrt{2}\simeq(\hat{b}_{y}\mp i\langle\hat{b}_{z}\rangle)/\sqrt{2}$, where $\langle\hat{b}_{z}\rangle=\sqrt{P_{0}/(\hbar\Omega_{0})}e^{-i\Omega_{0}t}$ with $P_{0}$ ($\Omega_{0}$) the power (angular frequency) of the incident light [44]. Because in the present case $\hat{b}$ and $\hat{b}^{\dagger}$ represent itinerant photon operators whose dimension is $\mathrm{time}]^{-1/2}$, the interaction Hamiltonian (16) needs to be integrated over the interaction time [30].
In order for the interaction time $\tau=d_{\mathrm{AFI}}/c$ with $d_{\mathrm{AFI}}$ the thickness of the antiferromagnet to be much shorter than the time scale of the antiferromagnetic dynamics (i.e., the antiferromagnetic resonance frequency) $1/\omega_{\mu}\sim 10^{-12}\,\mathrm{s}$, the thickness $d_{\mathrm{AFI}}$ needs to be of, or shorter than, $\mathcal{O}(1\,\mathrm{\mu m})$. When this condition is satisfied, the integrand, i.e., the operators in Eq. (16) can be regarded as constant during the interaction, enabling the integration over the interaction time as $\int_{0}^{\tau}dt=\tau$. Then, setting $\hat{b}_{y}\equiv\hat{b}_{\mathrm{in}}$, we arrive at the interaction Hamiltonian of the form
@@
$\times\left(\hat{b}_{\mathrm{in}}e^{i\Omega_{0}t}-\hat{b}^{\dagger}_{\mathrm{in}}e^{-i\Omega_{0}t}\right),$ (17)
@@
with $\phi_{\mathrm{F}}=\theta_{\mathrm{F}}d_{\mathrm{AFI}}$ being the total Faraday rotation angle per sublattice of the antiferromagnet (which should not be confused with $\theta_{\mathrm{F}}$).
@@ V Microwave-to-Optical Quantum Transduction
where $H_{0}=\hbar\omega_{\mathrm{e}}\hat{a}^{\dagger}\hat{a}-\hbar\delta\omega_{\mathrm{o}}\hat{b}^{\dagger}\hat{b}+\hbar\omega_{\alpha}\hat{m}_{\alpha}^{\dagger}\hat{m}_{\alpha}+\hbar\omega_{\beta}\hat{m}_{\beta}^{\dagger}\hat{m}_{\beta}$ for systems with microwave and optical cavities, while $H_{0}=\hbar\omega_{\mathrm{e}}\hat{a}^{\dagger}\hat{a}+\hbar\omega_{\alpha}\hat{m}_{\alpha}^{\dagger}\hat{m}_{\alpha}+\hbar\omega_{\beta}\hat{m}_{\beta}^{\dagger}\hat{m}_{\beta}$ for systems with a microwave cavity but without an optical cavity. Here, $\delta\omega_{\mathrm{o}}\equiv\omega_{\mathrm{P}}-\omega_{\mathrm{o}}$ is the detuning of the optical cavity frequency from the pump frequency [29].
@@ V.1 General consideration
where $S=I_{4\times 4}-B^{T}[-i\omega I_{4\times 4}+A]^{-1}B$ with $I_{4\times 4}$ being the $4\times 4$ identity matrix and we have used the Fourier transform defined by $\vec{c}(t)=\int dt/(2\pi)\,e^{-i\omega t}\vec{c}(\omega)$. Finally, the microwave-to-optical transduction efficiency $\eta$ is defined by
@@ V.2 The case with an optical cavity
$\dot{\hat{b}}=i\delta\omega_{\mathrm{o}}\hat{b}-i\sum_{\mu}\zeta_{\mu}\hat{m}_{\mu}-\frac{\kappa_{\mathrm{o}}}{2}\hat{b}-\sqrt{\kappa_{\mathrm{o,e}}}\hat{b}_{\mathrm{in}},$ (26)
@@
$A$ $=\begin{bmatrix}i\omega_{\mathrm{e}}+\kappa_{\mathrm{e}}/2&ig_{\alpha}&ig_{\beta}&0\\ ig_{\alpha}&i\omega_{\alpha}+\gamma_{\alpha}/2&0&i\zeta_{\alpha}\\ ig_{\beta}&0&i\omega_{\beta}+\gamma_{\beta}/2&i\zeta_{\beta}\\ 0&i\zeta_{\alpha}&i\zeta_{\beta}&-i\delta\omega_{\mathrm{o}}+\kappa_{\mathrm{o}}/2\end{bmatrix}$ (28)
@@
respectively. Finally, introducing the susceptibilities $\chi_{\mathrm{e}}=[-i(\omega-\omega_{\mathrm{e}})+\kappa_{\mathrm{e}}/2]^{-1}$, $\chi_{\mu}=[-i(\omega-\omega_{\mu})+\gamma_{\mu}/2]^{-1}$, and $\chi_{\mathrm{o}}=[-i(\omega+\delta\omega_{\mathrm{o}})+\kappa_{\mathrm{o}}/2]^{-1}$, the transduction efficiency $\eta$ of the microwave-to-optical quantum transduction is obtained from Eq. (23) as